The vertices of a right triangle are (6,5), (2,1), and (8,1). What are the coordinates of the midpoint of the hypotenuse?
Answer Choices:
A) (6,3)
B) (4,3)
C) (5,1)
D) (8,3)

Answer by checkley71(8405)

The hypotenuse of this triangle is the line between 2,1 & 8,1 with a length of
8-2=6 so the midpoint would be 6/2+2=3+2=5 for the x cordinate & 1 for the y coordinate or C)(5,1) is the midpoint.